OpenStudy (mathsucks321):

Write the standard form of the equation of the line that passes through the given points. After your write the equation, explain your steps to earn full points. through: (3, -3) and (-5, 4)

OpenStudy (mathsucks321):

Please help T_T

OpenStudy (gunsodo):

Okay, typing...

OpenStudy (gunsodo):

First find slope..

OpenStudy (gunsodo):

slope = difference of y/ difference of x so slope will be (-3-4)/(3-(-5)) = -7/8

OpenStudy (gunsodo):

Replace slope into the equation, y = mx + b ---> y = -7/8x+b

OpenStudy (gunsodo):

The left is for you to work it out!! Find b by using the coordinate, a simple equation solving needed

OpenStudy (mathsucks321):

I'm not sure please help walk me through this! @gunsodo

OpenStudy (gunsodo):

So, continue lol

OpenStudy (mathsucks321):

I don't know how I have trouble with these

OpenStudy (gunsodo):

After we've got y=-7/8 x+b we input (3,-3) into the equation we've just gotten

OpenStudy (mathsucks321):

How to I input it?

OpenStudy (gunsodo):

So, we'll get -3 = -7/8 (3) + b And b = -3/8

OpenStudy (gunsodo):

Hope my calculation is correct, please recheck. I am now totally sleepy haha

OpenStudy (mathsucks321):

Ty!

OpenStudy (gunsodo):

Sure:)